OpenStudy (anonymous):

Okay so how do you find the solution set to the system of equations: 4x^2 + 9y^2 = 72 and 2x - y = 4??

OpenStudy (anonymous):

2x-y =4 y = 2x-4 put this in eqn 1 4x^2+9(2x-4)^2=72 4x^2+9(4x^2+16-16x)=72 40x^2-144x+72=0 solving using quadratic equations x=3 and x=0.6 corresponding y=2 and y=-2.8

OpenStudy (anonymous):

Okay thank You! I'm glad you explained it all out haha I was so lost

OpenStudy (anonymous):

I do have a question though shouldn't the 16x in the third step be a 8x instead?